cLaval produces lamps and home lighting fixtures. Its most popular product is a brushed aluminum desk lamp. This lamp is made from components shaped in the fabricating department and assembled in the assembly department. Information related to the 35,000 desk lamps produced annually follows. Direct materials $ 255,000 Direct labor Fabricating department (9,000 DLH × $33 per DLH) $ 297,000 Assembly department (16,000 DLH × $22 per DLH) $ 352,000 Machine hours Fabricating department 14,800 MH Assembly department 20,100 MH Expected overhead cost and related data for the two production departments follow. Fabricating Assembly Direct labor hours 200,000 DLH 350,000 DLH Machine hours 157,000 MH 135,000 MH Overhead cost $ 410,000 $ 420,000 Required 1. Determine the plantwide overhead rate for Laval using direct labor hours as a base. 2. Determine the total manufacturing cost per unit for the aluminum desk lamp using the plantwide overhead rate. 3. Compute departmental overhead rates based on machine hours in the fabricating department and direct labor hours in the assembly department. 4. Use departmental overhead rates from requirement 3 to determine the total manufacturing cost per unit for the aluminum desk lamps.
Answer #1

Requirement #1:

Total direct labor hours for both departments:

= 200,000 + 350,000

= 550000 DLH

Total overhead cost = $ 410,000 + $ 420,000

= $830,000

Plantwide overhead rate on the basis of direct labor hours :

= $830000 / 550000

= $1.509 per DLH

Requirement#2:

Computation of manufacturing cost per unit - Plantwide overhead rate
Particulars Amount
Total Manufacturing cost:
Direct material $255,000
Direct Labor:
Fabricating department $ 297,000
Assembly department $ 352,000
Applied overhead (25000* x $1.509) $37,725
*25000=9000+16000
Total manufacturing cost $941,725
Total desk lamp produced 35,000
Manufacturing cost per unit(941725/35000) $26.906

Requirement#3:

Computation of departmental overhead rate
Particulars Fabricating Assembly
Overhead cost of deparment $410000.00 $420000.00
Allocation base MH DLH
Total machine hours 157000
Total direct labor hours 350000
Overhead rate of department

$2.611

$1.2

Requirement#4:

Computation of manufacturing cost per unit - Departmental overhead rate
Particulars Amount
Total Manufacturing cost:
Direct material $255000
Direct Labor:
Fabricating department $297000
Assembly department $352000
Applied overhead:
Fabricating department (14800*$2.611) $38,642.8
Assembly department (16000*$1.20) $19,200
Total manufacturing cost $961,842.8
Total desk lamp produced 35,000
Manufacturing cost per unit $27.48

ANY DOUBTS OR CORRECTIONS?

FEEL FREE TO COMMENT BELOW